# Enhancing Alzheimer’s Diagnosis: Leveraging Anatomical Landmarks in Graph Convolutional Neural Networks on Tetrahedral Meshes

### v1

@@ Abstract
Keywords:
Geometric deep learning Anatomical landmarking Tetrahedral meshes Cortex Alzheimer’s disease
@@ 1 Introduction
Alzheimer’s disease (AD) has been recognized as a major health problem since the past century, affecting patient cognition and causing dementia [28, 9]. Although there is still no cure for AD, several treatments have been developed to alleviate clinical symptoms and slow the disease progression [4, 22]. As a result, techniques for early diagnosis of Alzheimer’s disease receive attention among researchers and clinicians. Among brain imaging techniques, magnetic resonance imaging (MRI) has emerged as a leading and cost-effective method for detecting brain structural changes associated with AD, such as atrophy patterns over time [7]. Compared to more specialized techniques like positron emission tomography (PET), MRI is less invasive, less costly, and more accessible [22, 3].
While MRI images can be represented by 3-D voxels, the limited grid resolution prevents detailed study of arbitrarily small regions [12]. Volumetric mesh data structures offer comprehensive insights by modeling both surface and interior aspects of substances. As the most typical and elemental type of volumetric mesh, tetrahedral meshes have been widely used in scientific modeling such as finite element analysis and physical simulations [5]. Graph neural networks (GNNs) present a natural choice when analyzing graphs, which lack irregular grid-like structures as images. While algorithms based on graph convolution have developed for more commonly used surface meshes [17], convolutions on surface meshes suffer from the over-squashing nature of the message-passing neural networks, potentially aggregating unnecessary or redundant information from intermediate nodes when interacting with long-range nodes. This drawback can naturally be resolved by using volumetric mesh representation, as more detours are available in pathfinding. One notable method for tetrahedral mesh analysis is TetCNN, which was proposed as an improved graph neural networks for tetrahedral mesh data analysis [11] by replacing graph Laplacian with volumetric Laplace Beltrami operator (volumetric LBO) [33, 32].
Attention mechanism and transformer architectures are being used ubiquitously in deep learning applications [30, 8, 31]. Fundamental to the attention mechanism is a set of vectors associated with the input data, namely query vectors, key vectors, and value vectors. The outputs of each attention layer consist of the weighted sum of the value vectors, where the weights are determined by attention scores representing the similarity between the query and key vectors. Transformer architectures have proved their superiority over traditional CNNs in various tasks, such as machine translation, text generation, image classification and graph classification. However, these methods are not directly suitable for tetrahedral meshes with varying sizes and topologies, necessitating a novel transformer-based framework tailored to this structure and scalable to very large volumetric meshes.
@@ 2.1 Notation
where $N(i)$ are vertices adjacent to $v_{i}$, $d_{i}$ is the volume of adjacent tetrahedra, and $k_{i,j}$ depends on edge lengths and dihedral angles. The stiffness matrix $A=W-K$, with $W=\text{diag}(w_{1},\dots,w_{n})$ and $w_{i}=\sum_{j\in N(i)}k_{i,j}$. The lumped discrete LBO is $L_{tet}=D^{-1}A$ where $D=\text{diag}(d_{1},\dots,d_{n})$.
@@ 2.4 Node and Patch-Level Embeddings
In this equation $p$ is an arbitrary patch on the mesh, and $|p|$ shows the number of nodes in the patch. $X$ shows the patch embedding. With these embeddings, the tokens are now prepared. One approach is to add positional encodings and input them into a conventional transformer. However, due to the computational cost of full self-attention and the opportunity to leverage geometric information, a graph-based approach for calculating attention is more suitable. Using the super nodes as the representative of the tokens, we build a radius graph based o…….
@@ 2.5 Super Node Representation
$\mathcal{S}=\{s_{1},s_{2},\ldots,s_{N}\},\quad s_{i}=(p_{i},X_{i}),\quad p_{i}\in\mathbb{R}^{3},\;X_{i}\in\mathbb{R}^{d}.$
@@ 3.2 Data Preprocessing
The collected MRI images were first processed by Freesurfer to reconstruct the pial and white surfaces. We then generated the tetrahedral meshes using the TetGen pipeline [16], resulting in tetrahedral meshes containing around 70k-100k vertices. Finally, 1000 anatomical landmarks were extracted from each tetrahedral mesh using a GP-based landmarking generation pipeline.
To evaluate our model performance on brain amyloid positivity prediction task, we first identified ground truth class labels by computing the Centiloid values from PET images. Specifically, the PET images were processed by PET Unified Pipeline (PUP), and the positivity threshold was set to $Centiloid>20$, as proposed and validated in previous studies [21]. In addition, we applied a two-threshold approach to classify individuals into low, medium and high risk of being amyloid positive  [14]. The two thresholds were 1.53 and 2.602, determined by 95% sensitivity and 95% specificity, respectively. Since the pTau-217 alone has already achieved significant success as a sole predictor of brain amyloid positivity (AUC=0.94 (95% CI 0.92–0.97)), but not in the medium group (as shown in our results), we focus our effort on the medium group in our study [1].
@@ 4.1 Alzheimer’s Disease Classification
LETetCNN 0.765 0.738 0.790
LETetCNN+pTau-217 0.820 0.841 0.803
Table 3: Amyloid positivity prediction results in the medium risk groups. While pTau-217 performs well in general [29, 15, 26], its performance dropped significantly in the median risk group. An integration of LETetCNN and pTau-217 showed enhanced accuracy, outperforming all other baseline classifiers.
@@ 4.2 Brain Amyloid Positivity Classification
We generated a total of 909 brain cortical tetrahedral meshes from structural MR images. All samples were from ADNI dataset with matching PET scans and pTau-217 measures. We followed the widely used protocol [21] to determine brain amyloid positivity by calculating Centiloid value for each PET image, and labeled any PET image with $Centiloid>20$ as amyloid positive. We employed the aforementioned two-threshold method to classify samples into low, medium and high groups, and primarily focus on the medium group, where pTau-217 alone has demonstrated limited classification power. The subset of the medium group contained 70 positive and 96 negative samples. Hippocampal volume was also included as a baseline classifier, given its well-documented correlation with AD-related dementia. However, as shown in Table 3, hippocampal volume yielded results close to random classification. In comparison, pTau-217 alone achieved 0.675 classification accuracy, with similar performance observed from other GNN-based methods. In contrast, our LETetCNN model outperformaned all baseline models by achieving an accuracy of 0.765. Notably, the performance was further enhanced after integrating pTau-217, reaching a classification accuracy of 0.862 with a balanced prediction (SEN=0.917, SPE=0.824). While further validations are needed, our current results suggest that our work could complement BBBM research, enhancing its accuracy through a cost-effective and less invasive approach.
@@ 4.3 Ablation Studies
We first investigated the impact of learned features extracted by a few TetCNN layers before they were passed to the transformer. To assess this, we removed these layers and directly feed the raw features into the transformer. In the second ablation stude, we evaluated whether integrating $p$Tau-217 with LETetCNN significantly enhanced performance. This analysis was crucial, as $p$Tau-217 itself was a highly predictive biomarker, and it was important to determine whether the features learned by our network provided additional complementary information. The results of both studies were summarized in Table 4. It demonstrated TetCNN learned important geometric features and the proposed framework further enhanced the performance of BBBMs for preclinical AD research.
@@ 4.4 Grad-CAM Visualization
Figure 3: Grad-CAM visualization highlighting regions of interest that align with Alzheimer’s Disease (AD) pathology. Both results from AD classification task (A-C) and brain amyloid positivity prediction task (D-F) exhibited strong activations in the temporal gyrus and parts of the inferior temporal lobe, consistent with early cortical atrophy in AD. Mild activation was also observed in superior frontal cortex, which was related to later-stage disease progression. These results closely align with neuroimaging studies identifying key biomarkers for AD pathology.
@@ 5 Conclusion and Future Work
In this work, we proposed a novel transformer-based framework tailored specifically for tetrahedral meshes with varying sizes and topologies. To the best of our knowledge, this is the first approach to extend transformer architectures to tetrahedral meshes, leveraging relative positional encodings and task-specific graph tokenization for efficient and geometry-aware representation learning. Unlike prior methods that rely on fixed-size re-meshing or hierarchical pooling, our approach dynamically adapts to the inherent variability of tetrahedral mesh structures. We evaluated our model on two distinct tasks: clinical classification of the clinical AD groups and prediction of brain amyloid positivity. We compared our model performance with baseline classifiers, including ChebyNet and GAT in AD classification, and hippocampal volume and blood-based biomarker pTau-217. In all experiments, our model outperformed all baseline methods, indicating both effectiveness and generalizability. In particular, our LETetCNN model can be empowered by leveraging the pTau-217 information. With the rapid advancement of BBBM studies in AD research [29, 15, 26], our work indicates that traditional brain imaging research will still be useful and relevant.
Compared to TetCNN, which utilizes localized filters defined on the graph Laplacian and static downsampling for feature aggregation, our method introduces several key advantages. First, the attention mechanism in our framework enables dynamic modeling of global relationships across the mesh, overcoming the locality limitation of TetCNN layers. While TetCNN layers effectively capture local features, their hierarchical pooling may result in loss of critical geometric details, particularly for meshes with irregular or anisotropic topologies. In contrast, our method retains finer global and local information through attention mechanisms that explicitly consider both feature similarity and geometric relationships.
@@
Empirical results validate the effectiveness of our approach. Our framework achieves superior performance in clinical group classification, with an average of 5% increase in accuracy compared to pure TetCNN layers. Visualizations of attention maps further highlight the ability of our model to focus on critical regions, demonstrating richer feature representations. Furthermore, ablation studies show the significant contributions of relative positional encodings and landmark-based tokenization to overall performance.
@@
In conclusion, the proposed method represents a significant step forward in adapting transformer architectures to tetrahedral meshes, providing a robust, scalable, and geometry-aware solution for complex volumetric data. Future work can explore multi-scale tokenization strategies, integration of additional anatomical priors, and further applications in domains requiring flexible representation learning for 3D structures.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Alzheimer’s disease (AD) has been recognized as a major health problem since the past century, affecting patient cognition and causing dementia [28, 9]. Although there is still no cure for AD, several treatments have been developed to alleviate clinical symptoms and slow the disease progression, necessitating the development of early AD diagnosis techniques [3, 22]. Among brain imaging techniques, magnetic resonance imaging (MRI) has emerged as a leading and cost-effective method for detecting brain structural changes associated with AD [7]. Compared to more specialized techniques like positron emission tomography (PET), MRI is less invasive, less costly, and more accessible [22, 4].
While MRI images can be represented by 3-D voxels, the limited grid resolution prevents detailed study of arbitrarily small regions [12]. Volumetric mesh data structures, such as tetrahedral meshes, offer comprehensive insights by modeling both surface and interior aspects of substances, and have been widely used in scientific modeling such as finite element analysis and physical simulations [5]. Graph neural networks (GNNs) present a natural choice when analyzing graphs, which lack irregular grid-like structures as images. While algorithms based on graph convolution have developed for more commonly used surface meshes [17], convolutions on surface meshes suffer from the over-squashing nature of the message-passing neural networks, potentially aggregating unnecessary or redundant information from intermediate nodes when interacting with long-range nodes. This drawback can naturally be resolved by using volumetric mesh representation, as more detours are available in pathfinding. One notable method for tetrahedral mesh analysis is TetCNN, which was proposed as an improved graph neural networks for tetrahedral mesh data analysis [11] by replacing graph Laplacian with volumetric Laplace Beltrami operator (volumetric LBO) [33, 32].
Attention mechanism and transformer architectures are being used ubiquitously in deep learning applications [30, 8, 31]. Fundamental to the attention mechanism is a set of vectors associated with the input data, namely query vectors, key vectors, and value vectors. The outputs of each attention layer consist of the weighted sum of the value vectors, where the weights are determined by attention scores representing the similarity between the query and key vectors. Transformer architectures have proved their superiority over traditional CNNs in various tasks, such as machine translation, text generation, image classification and graph classification. However, these methods are not directly suitable for tetrahedral meshes with varying sizes and topologies, necessitating a novel transformer-based framework tailored to this structure with strong scalability.
@@ 2.1 Notation
where $N(i)$ are vertices adjacent to $v_{i}$, $d_{i}$ is the volume of adjacent tetrahedra, and $k_{i,j}$ depends on edge lengths and dihedral angles. The stiffness matrix A=WKA=W-K, with $W=\text{diag}(w_{1},\dots,w_{n})$ and $w_{i}=\sum_{j\in N(i)}k_{i,j}$. The lumped discrete LBO is $L_{tet}=D^{-1}A$ where $D=\text{diag}(d_{1},\dots,d_{n})$.
@@ 2.4 Node and Patch-Level Embeddings
In this equation $p$ is an arbitrary patch on the mesh, and $|p|$ shows the number of nodes in the patch. $X$ shows the patch embedding. With these embeddings, the tokens are now prepared. One approach is to add positional encodings and input them into a conventional transformer. However, due to the computational cost of full self-attention and the opportunity to leverage geometric information, a graph-based approach for calculating attention is more suitable.
@@ 2.5 Super Node Representation
$\mathcal{S}=\{s_{1},s_{2},\ldots,s_{N}\},\quad s_{i}=(p_{i},e_{i}),\quad p_{i}\in\mathbb{R}^{3},\;e_{i}\in\mathbb{R}^{d}.$
@@ 2.6 Attention Mechanism
2.7 Integrating BBBMs
After extracting the graph features, we integrated blood-based biomarkers by concatenating the BBBM values with the feature vector in the first fully connected layer, resulting in $d+1$ dimension inputs.
@@ 3.2 Data Preprocessing
The collected MRI images were first processed by Freesurfer to reconstruct the pial and white surfaces. We then generated the tetrahedral meshes using the TetGen pipeline [16], resulting in tetrahedral meshes containing around 70k-100k vertices. Finally, 1000 anatomical landmarks were extracted from each tetrahedral mesh using a GP-based landmarking generation pipeline. The ground truth class labels were identified by computing the Centiloid values from PET images. Specifically, the PET images were processed by PET Unified Pipeline (PUP), and the positivity threshold was set to $Centiloid>20$, as proposed and validated in previous studies [21]. In addition, we applied a two-threshold approach to classify individuals into low, medium and high risk of being amyloid positive  [14]. The two thresholds were 1.53 and 2.602, determined by 95% sensitivity and 95% specificity, respectively. Since the pTau-217 alone has already achieved significant success as a sole predictor of brain amyloid positivity (AUC=0.94 (95% CI 0.92–0.97)), but not in the medium group (as shown in our results), we focus our effort on the medium group in our study [1].
@@ 4.1 Alzheimer’s Disease Classification
LETetCNN 0.758 0.785 0.733
LETetCNN+pTau-217 0.798 0.785 0.811
Table 3: Amyloid positivity prediction results in the medium risk groups. An integration of LETetCNN and pTau-217 showed enhanced accuracy, outperforming all other baseline classifiers.
@@ 4.2 Brain Amyloid Positivity Classification
We generated a total of 909 brain cortical tetrahedral meshes from structural MR images. All samples were from ADNI dataset with matching PET scans and pTau-217 measures. We followed the widely used protocol [21] to determine brain amyloid positivity by calculating Centiloid value for each PET image, and labeled any PET image with $Centiloid>20$ as amyloid positive. We employed the aforementioned two-threshold method to classify samples into low, medium and high groups, and primarily focus on the medium group, where pTau-217 alone has demonstrated limited classification power. The subset of the medium group contained 70 positive and 96 negative samples. Hippocampal volume was also included as a baseline classifier, given its well-documented correlation with AD-related dementia. However, as shown in Table 3, hippocampal volume yielded results close to random classification. In comparison, pTau-217 alone achieved 0.675 classification accuracy, with similar performance observed from other GNN-based methods. In contrast, our LETetCNN model outperformaned all baseline models by achieving an accuracy of 0.758. Notably, the performance was further enhanced after integrating pTau-217, reaching a classification accuracy of 0.798 with a balanced prediction (SEN=0.785, SPE=0.811). While further validations are needed, our current results suggest that our work could complement BBBM research, enhancing its accuracy through a cost-effective and less invasive approach.
@@ 4.3 Ablation Studies
We first investigated the impact of learned features extracted by a few TetCNN layers before they were passed to the transformer. To assess this, we removed these layers and directly feed the raw features into the transformer. In the second ablation stude, we evaluated whether integrating ppTau-217 with LETetCNN significantly enhanced performance. This analysis was crucial, as ppTau-217 itself was a highly predictive biomarker, and it was important to determine whether the features learned by our network provided additional complementary information. The results of both studies were summarized in Table 4. It demonstrated TetCNN learned important geometric features and the proposed framework further enhanced the performance of BBBMs for preclinical AD research.
@@ 4.4 Grad-CAM Visualization
Figure 3: Grad-CAM visualization highlighting regions of interest that align with Alzheimer’s Disease (AD) pathology. Both results from AD classification task (A-C) and brain amyloid positivity prediction task (D-F) exhibited strong activations in the temporal gyrus and parts of the inferior temporal lobe. Mild activation was also observed in superior frontal cortex, which was related to later-stage disease progression. These results closely align with neuroimaging studies identifying key biomarkers for AD pathology.
@@ 5 Conclusion and Future Work
In this work, we proposed a novel transformer-based framework tailored specifically for tetrahedral meshes with varying sizes and topologies. To the best of our knowledge, this is the first approach to extend transformer architectures to tetrahedral meshes, leveraging relative positional encodings and task-specific graph tokenization for efficient and geometry-aware representation learning. Unlike prior methods that rely on fixed-size re-meshing or hierarchical pooling, our approach dynamically adapts to the inherent variability of tetrahedral mesh structures. We evaluated our model on two distinct tasks: clinical AD classification and brain amyloid positivity prediction. We compared our model with baseline classifiers, including ChebyNet and GAT in AD classification, and hippocampal volume and pTau-217 in amyloid positivity prediction. In both experiments, our model outperformed all baseline methods, indicating both effectiveness and generalizability. In particular, our LETetCNN model can be empowered by leveraging the pTau-217 information. With the rapid advancement of BBBM studies in AD research [29, 15, 26], our work indicates that traditional brain imaging research will still be useful and relevant.
Compared to TetCNN, which utilizes localized filters defined on the volumetric LBO our method introduces several key advantages. First, the attention mechanism in our framework enables dynamic modeling of global relationships across the mesh, overcoming the locality limitation of TetCNN. While TetCNN layers effectively capture local features, their hierarchical pooling may result in loss of critical geometric details, particularly for meshes with irregular or anisotropic topologies. In contrast, our method retains finer global and local information through attention mechanisms that explicitly consider both feature similarity and geometric relationships.
@@
Empirical results validate the effectiveness of our approach. Our framework achieves superior performance in clinical group classification, with an average of 5% increase in accuracy compared to TetCNN. Visualizations of attention maps further highlight the ability of our model to focus on critical regions, demonstrating richer feature representations. Furthermore, ablation studies show the significant contributions of relative positional encodings and landmark-based tokenization to overall performance.